OpenStudy (anonymous):

State whether the given measurements determine zero, one, or two triangles. B = 86, b = 21, c = 18. Is it two triangles?

OpenStudy (yanasidlinskiy):

\(21\div \sin c =18\div \sin \left( 20 \right)\) first step

OpenStudy (anonymous):

1

OpenStudy (yanasidlinskiy):

then cross multiply the 2 proportions giving you.... 21sinc=18sin20

OpenStudy (yanasidlinskiy):

It is not one.

OpenStudy (anonymous):

it's one

OpenStudy (yanasidlinskiy):

then divide both sides by 20 and you get 0.293 since the value is less than one the answer is 2 triangles